### v1

@@ 6 Conclusions
There are some limitiations in our theoretical analysis. We retained the assumption Eq. 2 from Marin et al. [15]. Although we used a moment-matching criterion in our experiments, which satisfies this assumption, there are some other criterion under which this assumption may not hold. However, it is important to note that this assumption is only used to design the algorithm for optimizing the parameters. Within our error analysis framework, we only need to know the parameter error, and the assumption itself is not required. Therefore, if we introduce other stochastic optimization algorithms that do not require this assumption to optimize the parameters, we might be able to weaken this assumption. In our experiments, we proposed a self-normalized MAMIS algorithm but did not conduct an error analysis for it. We leave these issues for future work.
@@
implying that $3^{1-q}\left|\widehat{I}_{n}(f)-\mathbb{E}[f(X)]\right|^{q}$ is not greater than
$\big{|}\widehat{I}_{n}(f)-\widehat{I}_{n}(f\circ P_{R})\big{|}^{q}+\big{|}\widehat{I}_{n}(f\circ P_{R})-\mathbb{E}\left[f\circ P_{R}(X)\right]\big{|}^{q}+{\big{|}\mathbb{E}\left[f\circ P_{R}(X)\right]-\mathbb{E}[f(X)]\big{|}}^{q}.$
Therefore, in order to analyze the total error $\mathbb{E}\left|\widehat{I}_{n}(f)-\mathbb{E}[f(X)]\right|^{q}$, we need to analyze three parts, $\mathbb{E}\big{|}\widehat{I}_{n}(f)-\widehat{I}_{n}(f\circ P_{R})\big{|}^{q}$, $\mathbb{E}\big{|}\widehat{I}_{n}(f\circ P_{R})-\mathbb{E}\left[f\circ P_{R}(X)\right]\big{|}^{q}$ and ${\big{|}\mathbb{E}\left[f\circ P_{R}(X)\right]-\mathbb{E}[f(X)]\big{|}}^{q}$.
The first and third parts are related to the following lemma.
@@
For any $f\in G_{e}(M,B,k)$ with $20<k<$ 2, the projection error $\mathbb{E}\left|f(X)-f\circ P_{R}(X)\right|^{q}$ is not greater than
$2^{q}(\frac{\pi}{2})^{\frac{d}{2}-1}B^{q}\exp\left\{\frac{2-k}{2}qM(2qkM)^{\frac{k}{2-k}}\right\}(d+1)!!(R-1)^{d-2}\exp\left\{-\frac{1}{4}(R-1)^{2}\right\}.$
where the projection radius $R>\sqrt{2}+1$.
@@
(27) $\mathbb{E}\left|f(X)-f\circ P_{R}(X)\right|^{q}\leq(2\pi)^{-\frac{d}{2}}\int_{|x|\geqslant R-1}\left(|f(x)|+\left|f\circ P_{R}(x)\right|\right)^{q}e^{-|x|^{2}/2}\mathrm{d}x.$
@@
then Eq. 28 is not greater than
@@
Note that, for the first part, we have
$\mathbb{E}\left|\widehat{I}_{n}(f)-\widehat{I}_{n}(f\circ P_{R})\right|^{q}=$ $\mathbb{E}\left[\left|\frac{1}{n}\sum_{j=1}^{n}f\circ\Phi^{-1}\left(y_{j}\right)-\frac{1}{n}\sum_{j=1}^{n}f\circ P_{R}\circ\Phi^{-1}\left(y_{j}\right)\right|^{q}\right],$
@@
where we use each $\Phi^{-1}(y_{j})\sim N(\bm{0},I_{d})$ in Eq. 29.
For the third part $\left|\mathbb{E}\left[f(X)-f\circ P_{R}(X)\right]\right|^{q}$, by Jensen’s inequality, we have
@@
$\leq$ $C_{1}(R-1)^{d-2}\exp\left\{-\frac{1}{4}(R-1)^{2}\right\}.$
Therefore, using Lemma A.1, we obtain the first and third part both can be bounded by
$2^{q}(\frac{\pi}{2})^{\frac{d}{2}-1}B^{q}\exp\left\{\frac{2-k}{2}qM(2qkM)^{\frac{k}{2-k}}\right\}(d+1)!!(R-1)^{d-2}\exp\left\{-\frac{1}{4}(R-1)^{2}\right\}.$
For the remaining second part $\mathbb{E}\left|\widehat{I}_{n}(f\circ P_{R})-E[f\circ P_{R}(X)]\right|^{q}$, we obtain the following result.
@@
Combining the results of the previous three errors, for any $f\in G_{e}\left(M,B,k\right)$ with $20<k<$ 2 and any $R>\sqrt{2}+1$, the total error $\mathbb{E}\left|\widehat{I}_{n}(f)-E[f(X)]\right|^{q}$ can be bounded by
@@
where the constant $C_{1}$ is equal to
$3^{q-1}2^{q+1}(\frac{\pi}{2})^{\frac{d}{2}-1}B^{q}\exp\left\{\frac{2-k}{2}qM(2qkM)^{\frac{k}{2-k}}\right\}(d+1)!!$!
and $C_{2}$ is equal to $3^{q-1}2^{2qd}B^{q}C.$
@@
where the constant in $\mathcal{O}$ depends on $M,B,k,d,q$.
Acknowledgments

### v2

@@ 6 Conclusions
In our theoretical analysis, we retained the assumption Eq. 2 from Marin et al. [15]. Although we used a moment-matching criterion in our experiments, which satisfies this assumption, there are some other criterion under which this assumption may not hold. However, it is important to note that this assumption is only used to design the algorithm for optimizing the parameters. Within our error analysis framework, we only need to know the parameter error, and the assumption itself is not required. Therefore, if we introduce other stochastic optimization algorithms that do not require this assumption to optimize the parameters, we might be able to weaken this assumption. In our experiments, we proposed a self-normalized MAMIS algorithm but did not conduct an error analysis for it. We leave these issues for future work.
@@
implying that
$3^{1-q}\left|\widehat{I}_{n}(f)-\mathbb{E}[f(X)]\right|^{q}\leq$ $\big{|}\widehat{I}_{n}(f)-\widehat{I}_{n}(f\circ P_{R})\big{|}^{q}+\big{|}\widehat{I}_{n}(f\circ P_{R})-\mathbb{E}\left[f\circ P_{R}(X)\right]\big{|}^{q}$
$+{\big{|}\mathbb{E}\left[f\circ P_{R}(X)\right]-\mathbb{E}[f(X)]\big{|}}^{q}.$
Taking the expectation on both sides of this inequality, we obtain
$3^{1-q}\mathbb{E}\left|\widehat{I}_{n}(f)-\mathbb{E}[f(X)]\right|^{q}\leq$ $\mathbb{E}\big{|}\widehat{I}_{n}(f)-\widehat{I}_{n}(f\circ P_{R})\big{|}^{q}+\mathbb{E}\big{|}\widehat{I}_{n}(f\circ P_{R})-\mathbb{E}\left[f\circ P_{R}(X)\right]\big{|}^{q}$
$+{\big{|}\mathbb{E}\left[f\circ P_{R}(X)\right]-\mathbb{E}[f(X)]\big{|}}^{q}.$
(27) =:absent:=:= : $E_{1}+E_{2}+E_{3}$
$E_{1}$ and $E_{3}$ are related to the following lemma.
@@
For any $f\in G_{e}(M,B,k)$ with $20<k<$ 2, we have
$\mathbb{E}\left|f(X)-f\circ P_{R}(X)\right|^{q}\leq C_{0}(R-1)^{d-2}\exp\left\{-\frac{1}{4}(R-1)^{2}\right\},$
where $C_{0}=2^{q}(\frac{\pi}{2})^{\frac{d}{2}-1}B^{q}\exp\left\{\frac{2-k}{2}qM(2qkM)^{\frac{k}{2-k}}\right\}(d+1)!!$! and the projection radius $R>\sqrt{2}+1$.
@@
(28) $\mathbb{E}\left|f(X)-f\circ P_{R}(X)\right|^{q}\leq(2\pi)^{-\frac{d}{2}}\int_{|\bm{x}|\geqslant R-1}\left(|f(\bm{x})|+\left|f\circ P_{R}(\bm{x})\right|\right)^{q}e^{-|\bm{x}|^{2}/2}\mathrm{d}\bm{x}.$
@@
then Eq. 29 can be bounded as follows.
$2^{q}B^{q}(\frac{\pi}{2})^{\frac{d}{2}-1}\int_{R-1}^{\infty}r^{d-1}\exp\left\{qMr^{k}\right\}e^{-r^{2}/2}\mathrm{d}r$
@@
Note that, for $E_{1}$ in Eq. 27, we have
$E_{1}=$ $\mathbb{E}\left|\widehat{I}_{n}(f)-\widehat{I}_{n}(f\circ P_{R})\right|^{q}$
$=$ $\mathbb{E}\left[\left|\frac{1}{n}\sum_{j=1}^{n}f\circ\Phi^{-1}\left(y_{j}\right)-\frac{1}{n}\sum_{j=1}^{n}f\circ P_{R}\circ\Phi^{-1}\left(y_{j}\right)\right|^{q}\right],$
@@
where we use each $\Phi^{-1}(y_{j})\sim N(\bm{0},I_{d})$ in Eq. 30. For $E_{3}$ in Eq. 27, by Jensen’s inequality, we have
@@
Therefore, using Lemma A.1, we obtain $E_{1}$ and $E_{3}$ both can be bounded by
$C_{0}(R-1)^{d-2}\exp\left\{-\frac{1}{4}(R-1)^{2}\right\}.$
For the remaining $E_{2}$, we obtain the following result.
@@
Using Eq. 27 and combining the results of the previous three errors $E_{1},E_{2}$ and $E_{3}$, for any $f\in G_{e}\left(M,B,k\right)$ with $20<k<$ 2 and any $R>\sqrt{2}+1$, we can bound $\mathbb{E}\left|\widehat{I}_{n}(f)-E[f(X)]\right|^{q}$ by
@@
where the constant $C_{1}=3^{q-1}2C_{0}$ and $C_{2}=3^{q-1}2^{2qd}B^{q}C.$
@@
where the constant in big-$\mathcal{O}$ bound depends on $M,B,k,d,q$.